OpenStudy (supersquidsisters):

answer check? What is the rate of change for f(x) = 7 sin x − 1 on the interval from x = 0 to x = pi over 2 ? (1 point) 14 over pi pi over 14 pi over 4 4 over pi my answer is pi over 14. thank you for any help!!

OpenStudy (sila1453):

I got 14 over pi :S

jimthompson5910 (jim_thompson5910):

It sounds like you have a step mixed up. You should end up at 14/pi f(x) = 7sin(x) - 1 f(0) = 7*sin(0)-1 f(0) = -1 f(x) = 7sin(x) - 1 f(pi/2) = 7sin(pi/2) - 1 f(pi/2) = 6 change in y = 6 - (-1) = 7 change in x = pi/2 - 0 = pi/2 slope = (change in y)/(change in x) = 7/(pi/2) = 14/pi

OpenStudy (supersquidsisters):

oh, ok!! thank you guys so much!! @Sila1453 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (sila1453):

You're welcome